LINE PARAMETERS OF CARBON DIOXIDE IN THE 4850 CM^{−1} REGION

\underline{D. CHRIS BENNER}, V. MALATHY DEVI, EMILY NUGENT, Department of Physics, College of William and Mary, Williamsburg, VA 23187-8795; KEEYOON SUNG, LINDA R. BROWN, CHARLES E. MILLER, ROBERT A. TOTH, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109, U.S.A..

The spectral region near 4850 cm^{−1} is used to monitor atmospheric carbon dioxide, but current accuracies of the line intensities and line shape coefficients do not permit carbon dioxide mixing ratios to be obtained to 1 ppm (about one part in 400). To improve the line parameters, we are remeasuring the prominent CO_{2} bands in this region specifically to characterize the non-Voigt effects of line mixing and speed dependence at room temperature.

The laboratory spectra of air- and self-broadened CO_{2} have been recorded at a variety of pressures, path lengths, mixing ratios and resolutions (0.005 to 0.01 cm^{−1}) with two different Fourier transform spectrometers (the McMath-Pierce FTS at Kitt Peak and a Bruker 125 HR FTS at JPL). The line parameters of some 2000 transitions are being derived by simultaneous multispectrum fitting^{a} using a few dozen spectra encompassing a 230 cm^{−1} wide spectral interval. The rovibrational constants for line positions and the band intensities and Herman-Wallis coefficients are being retrieved directly from the spectra, rather than floating positions and intensities individually.^{b} Self and foreign Lorentz widths and pressure shifts are being determined for the stronger bands while non-Voigt coefficients describing line mixing and speed dependence are being obtained for at least one of the strongest bands.^{c}

aD. Chris Benner, C. P. Rinsland, V. M. Devi, M. A. H. Smith, and D. A. Atkins, JQSRT 1995;53:705-21.

bV. M. Devi, D. Chris Benner, L. R. Brown, C. E. Miller, and R. A. Toth, J. Mol. Spectrosc. 2007;245:52-80.

cPart of the research described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration. Support for the work at William and Mary was provided by contracts with JPL.